Let B1, B2, B3 be pairwise exclusive, P(B1) = 1/8, P(B2) = 1/4, P(B3) = 5/8, P(A | B1) = 1, P(A | B2) = 1/3, P(A | B3) = 1/3. Calculate P(B2 | A).
Answer #1

Given that, B1, B2, B3 be pairwise exclusive, P(B1) = 1/8, P(B2) = 1/4, P(B3) = 5/8, P(A | B1) = 1, P(A | B2) = 1/3,

P(A | B3) = 1/3

We want to find, P(B2 | A)

P(AB2) P(B2) P(AB2)P(B2)1P(AB1) P(B1)1PAB3)P(B3)

P (B2|A)

P (B2|A) 1/121/8]+ [5/24]

2/2413/24+[5/24

P (B2|A) 2/243/241[5/24

P B2)+3+5)/24

PlB2IA) = (10124

(5/12)

PlB2|A) =-= 0.2

=> P(B2 | A) = 1/5 = 0.2